OpenStudy (anonymous):

Solve exactly in terms of ln. Be sure to check solutions: e^(2x) - e^x - 12 = 0

OpenStudy (anonymous):

okay. lets see. substitute e^x = y rewrite the expression in terms of y.

OpenStudy (anonymous):

did you try?

OpenStudy (anonymous):

y=e^(2x) - e^x - 12 ?

OpenStudy (anonymous):

no, lets say we substitute y wherever we see e^x. what do you get?

OpenStudy (anonymous):

y^2 - y - 12 = 0

OpenStudy (anonymous):

substitution basically

OpenStudy (anonymous):

yes, now do you see that it is a simple quadratic equation which can be factorized?

OpenStudy (anonymous):

which would be (y-4)(y+3)

OpenStudy (anonymous):

right. so what are the values of y?

OpenStudy (anonymous):

y= 4 or -3

OpenStudy (anonymous):

okay, now substitute back e^x for y and solve

OpenStudy (anonymous):

just to confirm?

OpenStudy (anonymous):

note that ln e^x = x

OpenStudy (anonymous):

so 4^2 - 4 - 12 = 0

OpenStudy (anonymous):

hmmm they both work

OpenStudy (anonymous):

yeah. it is a quadratic equation. it has 2 solutions.

OpenStudy (anonymous):

ln(4^2) - ln(4) - 12 ?

OpenStudy (anonymous):

I apologize still trying to grasp cpncept

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

y = 4 and y = -3 that means e^x = 4 and e^x = -3 ln e^x = ln 4 and ln e^x = ln -3 x = ln 4 and x = ln -3

OpenStudy (anonymous):

but we know that logrithms don't exist for negative numbers. so the answer is x = ln 4

OpenStudy (anonymous):

and thats it?